OpenStudy (anonymous):

(4y+6)^2=36 I'm helping a friend and she say's this is the answer below. answer is 0-3

OpenStudy (anonymous):

0, -3

OpenStudy (lalaly):

take the square root of both sides 4y+6= +/- 6 4y+6=6 4y=0 y=0 4y+6=-6 4y=-12 y=-3

OpenStudy (anonymous):

ummm easy thanks

OpenStudy (lalaly):

:)